OpenStudy (anonymous):

can someone solve this problem step by step for me please? 5 1/8 - 2 3/4 thanks!

OpenStudy (anonymous):

5 1/8 - 2 3/4 = 41/8 - 11/4 = 41/8 - 22/8 = 19/8 = 2 3/8 --> ans:)

OpenStudy (anonymous):

Convert the mixed fractions to improper fractions. 41/8 - 11/4 Multiply second fraction by 2 to get common denominator 41/8 - 22/8 19/8 as an improper fraction, 2 3/8 as a mixed fraction

OpenStudy (anonymous):

Nice explanation, EdgeOfTheEarth!

OpenStudy (anonymous):

Thanks so much Jay123 and EdgeOfTheEarth! We have been doing math homework for hours and have brain fog!